# MCGMark: An Encodable and Robust Online Watermark for LLM-Generated Malicious Code

### v1

@@ 6.1 Limitations
Another limitation is the impact of the watermark on the quality of LLM output. To address this, we design Algorithm 1 based on probabilistic outliers in the LLM’s vocabulary to ensure optimal token selection. However, the watermark inevitably impacts code quality. This occurs for two reasons: first, outliers themselves might have quality variations; second, the error correction bit of the watermark can influence the code output. Fortunately, the criteria for filtering outliers can be adjusted flexibly. A looser outlier filtering standard ensures more random model output, while a stricter standard guarantees higher output quality. Additionally, the introduction of the error correction bit reduces the watermark’s impact on model output by over $50\%$. Further compressing the length of the watermark, especially the error correction bit, can lead to even greater reductions in impact.
6.2 Threats to validity
6.2.1 Internal Validity
@@
In section 3, the construction of the MCGTest relied on manual analysis by participants, which introduces a certain degree of subjectivity to the results. To mitigate this threat, we use a close card sorting method for each aspect requiring manual analysis, involving at least two participants to ensure consistency in the results. We make significant efforts to mitigate this potential threat.
@@
Another internal validity concern in constructing MCGTest is our reliance on keyword matching for gathering code repositories and data, which may lead to incomplete data collection. However, the primary goal in creating MCGTest is to assemble a sufficient number of scenarios involving malicious code generation. This allows us to design, test, and inspire future developments in MCGMark. Our objective is not to capture every instance of malicious code but to collect representative cases that serve our research needs.
6.2.2 External Validity
@@
We test MCGMark exclusively on DeepSeek-Coder-6.7b because it has demonstrated high performance with relatively small parameters in various benchmarks [5]. However, the watermark embedding, detection, and robustness algorithms proposed in this paper are not specific to any single LLM. They are designed to be applicable to any LLM based on the transformer architecture [78]. Adapting MCGMark to other models involves adjusting the code element matching in Algorithm 2 to fit the model’s vocabulary. This adjustment is straightforward and does not present significant technical challenges.
Similarly, this paper adapts watermark patterns specifically for the Python language, as it is currently the most commonly used language among attackers [79]. Adapting the watermark patterns to other languages is not a technical challenge; it merely requires adjusting and matching the code elements of the target language with the seven patterns of code elements. It is important to note that since our watermark is embedded online, we do not have access to the entire code during embedding. Therefore, we cannot extend MCGMark’s adaptability based on the Abstract Syntax Tree (AST).
7 Discussion
@@ 7.1 Watermark Embedding Failure
Poor Model Outputs. Watermark embedding often fails in scenarios where model outputs are subpar, such as generating code that doesn’t adhere to syntax rules or mixes natural language with code. Online watermark embedding heavily relies on the generation capabilities of LLMs. Transitioning to a more powerful model is the most direct solution. Additionally, embedding considerations should include rollback strategies based on tokens generated by LLMs to safeguard model outputs.
7.2 Watermark Detection Failure
Tokenization Discrepancies. Tokenizers are not always reliable. Minor discrepancies between tokenization during code generation and code detection may lead to errors in watermark detection. Notably, if tokens affected by these discrepancies do not contain watermarks, automatic correction may occur during subsequent detection processes. Conducting systematic empirical studies on tokenizer errors could provide insights into addressing this issue.
@@ 7.2 Watermark Detection Failure
8 Related work
8.1 Traditional Code Watermark
Code watermarking involves directly adding a special identifier to the source code or during code execution to declare code ownership [80]. Code watermarking techniques can be broadly categorized into static and dynamic approaches [19]. Static watermarking embeds watermarks directly into the source code. For instance, Kim et al. [81] uses adaptive semantic-preserving transformations to embed watermarks, and Sun et al. [82] does so by changing the order of functions. However, static watermarks are relatively more susceptible to detection and removal [83]. Consequently, dynamic code watermarking techniques have seen rapid development recently. For example, LLWM [84] is a watermarking technique that uses LLVM and Clang to embed watermarks by compiling the code. Xmark [85] embeds watermarks by obfuscating the control flow based on the Collatz conjecture. However, dynamic watermarking techniques are not applicable to the code generation process of LLMs, as LLMs do not execute the generated code.
Difference. Therefore, current traditional code watermarking techniques are not suitable for LLM code generation. The watermarking techniques mentioned above differ significantly from the watermark proposed in this paper, which operates during the LLM code generation process. We need to design more innovative watermarks for the LLM code generation task.
8.2 LLM Watermark
The security of LLMs has recently gained wide attention, leading to efforts to use watermarking techniques for their protection. Regarding watermarking techniques for model security, the aim is to prevent model theft through methods such as distillation. For example, GINSW [86] injects private signals into the probability vectors of the decoding step for each target token to prevent model theft. TOSYN [18] replaces code training data samples with synthesized code to protect the model against distillation attacks. PLMmark [87] ensures model ownership by watermarks in the LLM training.
@@ 8.2 LLM Watermark
8.3 Watermarking for LLM Text Generation
Non-encodable Text Watermark. LLM watermarking techniques can be categorized into offline and online watermarking. In online watermarking, the watermarking process is synchronized with the LLM content generation process [88]. On the other hand, offline watermarking requires processing the generated text after the LLM has completed content generation [89, 90]. Offline watermarking is not directly related to the LLM itself. They rely more on rules and are easier for attackers to detect [20]. Kirchenbauer et al. [25] is a text watermarking technique that has received considerable attention. It involves partitioning the vocabulary of LLM and guiding the model to select words from a predefined vocabulary, thereby embedding watermarks into the generated text [88]. A follow-up work [88] investigates the reliability of this watermarking strategy. Subsequent works have extended this watermarking technique to privacy [91] and robustness [42].
@@ 8.3 Watermarking for LLM Text Generation
Encodable Text Watermark. Yoo et al. [43], also building on the work of Kirchenbauer et al. [25], achieved multi-bit watermark embedding by dividing the vocabulary into more sub-vocabularies. Wang et al. [92] analyzed how to embed more information into watermarks through vocabulary partitioning from a mathematical perspective. However, further subdividing the vocabulary may lead to a significant decline in the quality of the model’s output.
Difference. Therefore, the aforementioned works are only applicable to text-generation scenarios. Code, as a special type of text, has a relatively fixed structure and pattern. These watermarks are not designed for code generation scenarios and cannot address the issues proposed in this paper.
8.4 Watermarking for LLM Code Generation
As the issue of malicious developers using LLMs to generate malicious code becomes more recognized, some works have designed watermarking schemes specifically for LLM code generation. Li et al. [23] devised a set of code transformation rules to embed watermarks through post-processing. Yang et al. [24] designed an AST-based code transformation method to embed multi-bit watermarks, which is also a post-processing watermark. Post-processing watermarks are more susceptible to attackers discovering their rules and disrupting the watermark. Additionally, neither of these methods offers robust solutions tailored to the structure of the code. SWEET [21] is an online code watermark that extends the low-entropy scenarios of Kirchenbauer et al. [25] to improve code generation quality. However, SWEET can only achieve binary results and still does not address the issues proposed in this paper.
@@ 8.4 Watermarking for LLM Code Generation
9 Conclusion
In this work, we propose MCGMark, a robust and encodable watermarking technique for LLMs to counteract the growing trend of malicious code generation. MCGMark leverages code structural features to embed user identity information while LLM token output. To minimize the impact on LLM code generation quality, we introduce a watermark quality enhancement scheme for MCGMark based on probabilistic outliers. Furthermore, we design a watermark skipping algorithm based on code structural features and syntax rules, excluding easily removable elements such as comments, thus enhancing MCGMark robustness. We also develop MCGTest, a dataset of $406$ malicious code generation tasks, derived from a review of existing malicious code. And we evaluate the effectiveness and robustness of MCGMark on MCGTest.
We make MCGTest and MCGMark open-source to inspire further countermeasures against malicious code generation, addressing the growing problem of malicious developers using LLMs to create harmful code. Additionally, in future work, we plan to explore watermark compression methods to reduce interference with code generation quality. We also intend to develop new token-matching techniques to accommodate the complex tokens in different LLMs, thereby enhancing the defense capabilities of the watermark.

### v2

@@ 6. Limitations
Inevitable Impact on Code Quality. Another limitation is the impact of the watermark on the quality of LLM output. To address this, we design Algorithm 1 based on probabilistic outliers in the LLM’s vocabulary to ensure optimal token selection. However, the watermark inevitably impacts code quality. This occurs for two reasons: first, outliers themselves might have quality variations; second, the error correction bit of the watermark can influence the code output. Fortunately, the criteria for filtering outliers can be adjusted flexibly. A looser outlier filtering standard ensures more random model output, while a stricter standard guarantees higher output quality. Further compressing the length of the watermark, especially the error correction bit, can lead to even greater reductions in impact.
Unavoidable Overhead Introduction. A further limitation of MCGMark is the additional time overhead, a common issue for all watermarking schemes (Zhang et al., 2023a; Zhang and Koushanfar, 2024). Nevertheless, as shown in Section 5.6, its overall overhead remains comparable to the baselines. MCGMark also avoids loading extra models or databases during embedding and requires only the malicious code for detection, without relying on the LLM, external resources, or intermediate states. This design improves space efficiency and practicality for real-world SSP traceability. Future work will focus on optimizing hash computation and vocabulary partitioning to further reduce time and space complexity.
7. Threats to validity
7.1. Internal Validity
@@
Subjective Bias in Manual Analysis. In section 3, the construction of the MCGTest relied on manual analysis by participants, which introduces a certain degree of subjectivity to the results. To mitigate this threat, we use a close card sorting method for each aspect requiring manual analysis, involving at least two participants to ensure consistency in the results. We make significant efforts to mitigate this potential threat.
@@
Limitations of Data Collection Methods. Another internal validity concern in MCGTest is its reliance on keyword matching for data collection, which may be incomplete. However, the goal is to gather sufficient malicious code scenarios to design, test, and advance MCGMark. Rather than capturing every instance, we focus on representative cases that meet our research needs.
Scale Limitations of Evaluation Benchmarks. Additionally, we acknowledge that MCGMark was evaluated on the MCGTest, which may have scale limitations. However, with 406 tasks, MCGTest is already substantial. In comparison, widely used LLM benchmarks like OpenAI’s HumanEval (Chen et al., 2021) contain only 164 tasks. Moreover, as noted by CodeIP (Guan et al., 2024), benchmarks such as HumanEval and MBPP (Roziere et al., 2023) focus on simple problems with short generated code, making them unsuitable for assessing longer multi-bit watermark embedding.
7.2. External Validity
@@
Scope of Model Adaptability. We only evaluated MCGMark on three open-source LLMs, without extending the assessment to additional models. However, the watermark embedding, detection, and robustness algorithms in MCGMark are not tailored specifically to these evaluated LLMs. They are designed to be applicable to any LLM based on the transformer architecture (Kirchenbauer et al., 2023b). Adapting MCGMark to other models involves adjusting the code element matching in Algorithm 2 to fit the model’s vocabulary. This adjustment is straightforward and does not present significant technical barriers.
Generalizability Across Programming Languages. In this paper, watermark patterns are specifically tailored to Python due to its current dominance among attackers (Acarturk et al., 2021). However, extending these watermark patterns to other programming languages does not present substantial technical challenges. The process requires only minor modifications to map and match the target language’s code elements to our established watermark patterns. It is worth noting that because our watermark embedding process operates synchronously with code generation, we do not have access to the complete source code at embedding time. Thus, extending MCGMark’s adaptability using AST-based approaches is not feasible in the current design.
8. Discussion
@@ 8.1. Watermark Embedding
Poor Model Outputs. Watermark embedding often fails when model outputs are subpar, such as generating syntactically incorrect code or mixing natural language with code. Online watermarking relies heavily on the LLM’s generation capabilities, making a more powerful model the most direct solution. Additionally, embedding strategies should incorporate rollback mechanisms based on generated tokens to preserve output quality.
8.2. Watermark Detection
Tokenization Discrepancies. Tokenizers are not always reliable. Minor discrepancies between tokenization during code generation and code detection may lead to errors in watermark detection. Notably, if tokens affected by these discrepancies do not contain watermarks, automatic correction may occur during subsequent detection processes. Conducting systematic empirical studies on tokenizer errors could provide insights into addressing this issue. Additionally, due to its reliance on tokenizers, MCGMark is only applicable for detecting and tracing code generated by LLMs. It is not suitable for evaluating human-written code, as such code may not be accurately tokenized, potentially leading to unexpected results. Furthermore, the evaluation of human-written code falls outside the scope of this work.
@@ 8.2. Watermark Detection
9. Related work
9.1. Traditional Code Watermark
Code watermarking involves directly adding a special identifier to the source code or during code execution to declare code ownership (Kitagawa and Nishimaki, 2022). Code watermarking techniques can be broadly categorized into static and dynamic approaches (Li et al., 2023d). Static watermarking embeds watermarks directly into the source code. For instance, Kim et al. (Kim et al., 2023) uses adaptive semantic-preserving transformations to embed watermarks, and Sun et al. (Sun et al., 2023) does so by changing the order of functions. However, static watermarks are relatively more susceptible to detection and removal (Kang et al., 2021). Consequently, dynamic code watermarking techniques have seen rapid development recently. For example, LLWM (Novac et al., 2021) is a watermarking technique that uses LLVM and Clang to embed watermarks by compiling the code. Xmark (Ma et al., 2019) embeds watermarks by obfuscating the control flow based on the Collatz conjecture. However, dynamic watermarking techniques are not applicable to the code generation process of LLMs, as LLMs do not execute the generated code.
Difference. Therefore, current traditional code watermark are not suitable for LLM code generation. The watermarking techniques mentioned above differ significantly from the watermark proposed in this paper, which operates during the LLM code generation process. We need to design more innovative watermarks for the LLM code generation task.
9.2. LLM Watermark
The security of LLMs has recently attracted significant attention, leading to the adoption of watermarking techniques for protection. Watermarking aims to prevent model theft, particularly through distillation in model security. For instance, GINSW (Zhao et al., 2023) embeds private signals into probability vectors during decoding to deter theft. TOSYN (Li et al., 2023c) replaces training samples with synthesized code to defend against distillation attacks. PLMmark (Li et al., 2023a) embeds watermarks during LLM training to establish model ownership.
@@ 9.2. LLM Watermark
9.3. Watermarking for LLM Text Generation
Non-encodable Text Watermark. LLM watermarking techniques can be categorized into offline and online watermarking. In online watermarking, the watermarking process is synchronized with the LLM content generation process (Kirchenbauer et al., 2023c). On the other hand, offline watermarking requires processing the generated text after the LLM has completed content generation (Peng et al., 2023; Chang et al., 2024). Offline watermarking is not directly related to the LLM itself. They rely more on rules and are easier for attackers to detect (Liu et al., 2023b). Kirchenbauer et al. (Kirchenbauer et al., 2023a) is a text watermarking technique that has received considerable attention. It involves partitioning the vocabulary of LLM and guiding the model to select words from a predefined vocabulary, thereby embedding watermarks into the generated text (Kirchenbauer et al., 2023c). A follow-up work (Kirchenbauer et al., 2023c) investigates the reliability of this watermarking strategy. Subsequent works have extended this watermarking technique to privacy (Liu et al., 2024a) and robustness (Yoo et al., 2023).
@@ 9.3. Watermarking for LLM Text Generation
Encodable Text Watermark. Yoo et al. (Yoo et al., 2024), also building on the work of Kirchenbauer et al. (Kirchenbauer et al., 2023a), achieved multi-bit watermark embedding by dividing the vocabulary into more sub-vocabularies. Wang et al. (Wang et al., 2024) analyzed how to embed more information into watermarks through vocabulary partitioning from a mathematical perspective. However, further subdividing the vocabulary may lead to a significant decline in the quality of the model’s output. Boroujeny et al. (Boroujeny et al., 2024) proposed a multi-bit watermarking embedding scheme that does not alter the probability distribution of the vocabulary but instead controls the token selection offset of the LLM.
Difference. The aforementioned works are only applicable to text-generation scenarios. Code, as a special type of text, has a relatively fixed structure and pattern. These watermarks are not designed for code generation scenarios and cannot address the issues proposed in this paper.
9.4. Watermarking for LLM Code Generation
As the issue of malicious developers using LLMs to generate malicious code becomes more recognized, some works have designed watermarking schemes specifically for LLM code generation. Li et al. (Li et al., 2024b) devised a set of code transformation rules to embed watermarks through post-processing. Yang et al. (Yang et al., 2024a) designed an AST-based code transformation method to embed multi-bit watermarks, which is also a post-processing watermark. Post-processing watermarks are more susceptible to attackers discovering their rules and disrupting the watermark. Additionally, neither of these methods offers robust solutions tailored to the structure of the code. SWEET (Lee et al., 2024) is an online code watermark that extends the low-entropy scenarios of Kirchenbauer et al. (Kirchenbauer et al., 2023a) to improve code generation quality. However, SWEET can only achieve binary results and still does not address the issues proposed in this paper. CodeIP (Guan et al., 2024) designed a code-based multi-bit watermarking scheme, which restricts the sampling process of predicting the next token by training a type predictor.
@@ 9.4. Watermarking for LLM Code Generation
10. Conclusion
In this work, we propose MCGMark, a robust and encodable watermarking technique for LLMs to counteract the growing trend of malicious code generation. MCGMark embeds user identity information implicitly into the generated code by controlling the LLM’s token selection process. To improve watermark quality, MCGMark leverages probabilistic outliers in the vocabulary to optimize the quality of candidate tokens. Additionally, to enhance robustness, it uses code structure and syntax rules to skip easily modifiable elements such as comments, reducing the risk of watermark removal. Furthermore, we construct MCGTest, the first prompt dataset targeting LLM-generated malicious code. It consists of 406 tasks covering both real-world instances and potential risk scenarios. We conduct a comprehensive evaluation of MCGMark on this dataset, and the results demonstrate that it achieves strong performance in terms of embedding success rate, generation quality, robustness, and time efficiency. We have open-sourced both MCGMark and MCGTest to provide a reproducible foundation for the research community. In future work, we plan to explore watermark compression techniques to further reduce interference with code generation, and develop more general token-matching mechanisms to support complex token structures across different LLM architectures, thereby improving the generality and defense capability of the watermark.